OpenStudy (anonymous):

Assume g is a one-to-one function. If g(x)=x^2+10x with x>-5, find g^-1(10)

OpenStudy (anonymous):

you need to solve: \[10 = x^2+10x\] This is going to return 2 answers, and you pick the one that is greater than -5

OpenStudy (anonymous):

but it is g^-1

OpenStudy (anonymous):

thats right: \[g^{-1}(10) \] is the same as asking what value of x makes: \[g(x) = 10\] So we set g(x) = 10, and solve for x.

OpenStudy (anonymous):

i thought you have to do something special... i thought like this: lets call g^-1(x) = a g^-1(x)= a^2 +10a so it would turn out the same in this problem but in others it might be different right? i kinda forget even though i learned this like this year haha

OpenStudy (anonymous):

hmm...thats sorta right. if you see it like that, it would be: \[g^{-1}(10) = a\] \[g(x) = x^2+10x \Rightarrow 10 = a^2+10a\] and you would have to solve for a.

OpenStudy (anonymous):

yeah ok thanks for the clarification!